### v1

Abstract
This paper introduces and formalizes Noosemia, a novel cognitive-phenomenological phenomenon emerging from human interaction with generative AI systems, particularly those enabling dialogic or multimodal exchanges. We propose a multidisciplinary framework to explain how, under certain conditions, users attribute intentionality, agency, and even interiority to these systems—a process grounded not in physical resemblance, but in linguistic performance, epistemic opacity, and emergent technological complexity. By linking an LLM declination of meaning holism to our technical notion of the LLM Contextual Cognitive Field, we clarify how LLMs construct meaning relationally and how coherence and a simulacrum of agency arise at the human–AI interface. The analysis situates noosemia alongside pareidolia, animism, the intentional stance and the uncanny valley, distinguishing its unique characteristics. We also introduce a-noosemia to describe the phenomenological withdrawal of such projections. The paper concludes with reflections on the broader philosophical, epistemological, and social implications of noosemic dynamics and directions for future research.
Keywords: Generative AI, cognition, intentionality, phenomenology, semiotics, noosemia, human–AI interaction
Definition: Noosemia Noosemia (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological phenomenon in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemia thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemia refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
1 Introduction
“Technology is the procedure by which scientific man masters nature for the purpose of moulding his existence, delivering himself from want, and giving his environment the form that appeals to him. The appearance given to nature by human technology and how his technological procedure reacts upon man, that is, the manner in which his mode of work, work organisation, and shaping of the environment modify him himself, constitutes one of the ground-lines of history.”
— Karl Jaspers, [jaspers2010origin, p. 98]
The past few years have witnessed a huge and indubitable transformation in the landscape of artificial intelligence. The advent of generative models—most notably, large language models (LLMs) such as ChatGPT, Gemini, and Claude—has enabled a form of linguistic interaction never before seen in the history of technology. Indeed, we can safely say we are in the midst of a “cognitive revolution” [^1] That we are at the beginning of a “cognitive revolution” that will have a strong, widespread impact on knowledge and social systems is enough to analyze the statements [jaimungal2025hinton] of Nobel Prize winner Geoffrey Hinton—known as the godfather of AI— regarding generative models and the new era to come; see https://www.youtube.com/watch?v=b_DUft-BdIE. where “machines” are solving tasks and performing activities in domains once thought to be exclusively human. Thanks to their mastery of languages and other forms of expression, machines are effectively able to emulate previously unseen forms of thought and a ew kind of cognitive ergonomics. Unlike earlier automata, whose behavior was passive, transparent and predictably bounded, these new systems speak with a fluency, contextual sensitivity and apparent intentionality, provoking in the user not just curiosity or technical admiration, but a deeper sense of cognitive resonance and surprise.
A telling moment in this unfolding story comes from Sam Altman, CEO of OpenAI, who, in a widely discussed interview published on July 23, 2025 [altman2025podcast], recounted an episode that many have come to recognize as emblematic of the new relationship between humans and intelligent machines. After submitting a question to GPT-5—a question he himself admitted he could not fully understand—Altman was met with a response so apt, so clear and relevant, that he described himself as momentarily stunned. He sat back in his chair, struck by a feeling of astonishment, a recognition that the machine had accomplished in an instant what he could not do himself. It was, he said, a “weird feeling” (“oh man here it is moment” Altman said referring to the launch of GPT-5), a fleeting but powerful moment of displacement and wonder, one that echoes the experiences of countless others who have, in recent months, tested the boundaries of their own understanding in conversation with these new artificial interlocutors [^2] Specifically the Von interviewer after Altman story did this observation: “Yeah I think that’s I think that feeling right there that’s the feeling a lot of people kind of have like what’s going you know when does it happen what’s going to happen”; see https://www.youtube.com/watch?v=aYn8VKW6vXA. [altman2025podcast].
In another recent interview between Andrew Mayne and Sam Altman, the term “moment of AGI” explicitly appears in the conversation [openai2025podcast2]. Mayne observes that many users report experiencing such a moment when interacting with advanced agentic systems, such as Operator with o3333Operator is an AI agent developed by OpenAI that can autonomously control the user’s computer and browser to perform complex, multimodal tasks. The underlying model, “o3”, is OpenAI’s next-generation reasoning engine, designed to integrate external tools and operate on text, images, and live data in real time [openai2025o3, operator2025o3upgrade]., describing the impression of watching an AI “use a computer pretty well” [^4] See https://www.youtube.com/watch?v=DB9mjd-65gw%7D%7D. For Mayne, his own “moment of AGI” occurred with Deep Research, as he witnessed the system autonomously gather and synthesize information “better than I read before”, closely resembling a genuinely agentic investigative process. Altman, while noting that the effect was less pronounced for him personally, acknowledges that such moments are becoming widespread as AI capabilities evolve. These firsthand accounts help illustrate how users encounter a threshold of surprise and cognitive resonance while interacting with such smart AI systems [^5] We’re aware of the commercial logic behind this type of interview. However, this doesn’t negate the possibility of intense surprise during a specific interaction with a generative AI system..
What is it that lies at the heart of such experiences? The first encounter with a generative language model is (in the near future equipped with strong agentic capabilities), for many, a moment charged with surprise—a sensation that one’s own thoughts, intentions, or even doubts have been grasped, anticipated, or resolved by a machine. This “wow effect”, as it has come to be called, is not merely a function of technical proficiency, but of a much deeper alignment between linguistic performance and human cognitive expectation. The LLM does not simply return information; it emulates a form of sense-making that appears, at least in the moment, to rival or even surpass that of its human partner [bubeck2023, desantis2023apocalissi]—Fig 1 is a pictorial illustration of the “wow effect”.
Figure 1: Dialogic Flow and the “Wow Effect” in Human–LLM Interaction.
The illustration depicts the cycle of user input and LLM response, highlighting the emergence of the noosemic “wow effect”—a moment of surprise and cognitive resonance when the user perceives unexpected intelligence in the AI’s output (Image generated with OpenAI ChatGPT, image engine: DALL·E 3 in July 2025).
It is worth noting that the shock of this encounter is amplified by the “cognitive ergonomics” of language. Unlike earlier machines—which demanded users adapt to their logic and limitations—these models adapt themselves, flexibly and contextually, to human expression. This interaction, deeply embedded in natural language, fosters a sense of intimacy and mutual understanding, an illusion of shared meaning that far exceeds anything previously produced by mechanical or classical algorithmic processes [turkle2011alone, wang2024understanding]. As recent studies have shown, users frequently report feelings of connection, empathy, or even trust in their first interactions with LLMs, especially when the system provides responses that feel both novel and deeply relevant [cohn2024believing, wang2024understanding].
In attempting to explain this effect, one must take into account the technical details of neural architectures or the mathematical intricacies of transformer models. But there is, at play, an irreducible complexity—a layering of signs [eco1976theory, peirce1931collected], meanings, and expectations—rooted in both the architecture of the system and the architecture of human cognition. The user’s sense of surprise is not just a response to statistical fluency or syntactic correctness; it arises from a felt encounter with what seems like a mind at work, a co-constructor of meaning whose origins and intentions remain, in the end, hidden behind the black box of contemporary AI [lipton2018mythos, desantis2023apocalissi]. This leads to the era of “opaque machines”.
The opacity, or “explanatory gap”, is itself not new in the philosophy of mind. As Chalmers argues, the explanatory gap refers to the persistent difficulty in bridging the divide between the observable phenomena—such as linguistic sense or behavioral competence—and the underlying physical or computational mechanisms that give rise to them [chalmers1996conscious]. Yet, in the case of LLMs, it is compounded by the scale and hierarchy of the models, the astronomical size of their representational space due to the size of the context windows and their ability to generate not just plausible, but sometimes astonishingly creative, responses [brown2020language, wei2022emergent, bommasani2021opportunities]. The dialogue with an LLM, especially at first contact (or for a user with no technical experience), can take on the character of a magical interaction, reminiscent of the anthropological notion of “magic” explored in the past by Mauss, Lévy-Bruhl, and other classical theorists [mauss1906origins, lewybruhl1923primitive]. The user, unable to reconstruct the causal chains underlying the model’s output, finds herself in a position akin to that of a spectator at a prestidigitation show: she knows, rationally, that there is a “trick”, a mechanism, but is nonetheless enchanted by the effect [desantis2023apocalissi]. The spectator, who today has at least an average level of education, rationally “knows” that the predictor has a suite of magic tricks, yet in most cases remains surprised and in a state of doubt. Moreover, just as audiences willingly suspend disbelief and experience authentic emotional reactions when engaging with fictional characters in literature or film, despite being fully aware of their constructed nature, users of generative AI systems may find themselves empathizing with and attributing interiority to these artificial agents. Empirical studies have shown that such “narrative empathy” and the capacity for simulation are robust features of human cognition [mar2009exploring, green2000role, oatley2016fiction]. The boundaries between illusion and reality, between mind and mechanism, become fluid and negotiable.
On the other hand, Georg Simmel, in his analysis of modernity at the beginning of XX century, remarked on the ways in which knowledge is increasingly compartmentalized, encapsulated within technical systems whose workings become inaccessible to the layperson [simmel2004philosophiemoderne]. In the age of generative AI, this dynamic is raised to new heights; the black box is not just an engineering metaphor, but a lived phenomenological reality, one that reshapes our experience of agency, knowledge and self. As Altman’s anecdote illustrates, the user may experience both fascination and a kind of existential unease—wonder at the machine’s apparent intelligence and uncertainty about one’s own place in the emerging ecology of minds.
It is in response to this new condition that we introduce the concept of “Noosemia”, as a distinct specification of Dennett’s “intentional stance” [dennett1987intentional], that is adopting a strategy of interpreting an entity’s behavior by ascribing beliefs and desires to it. The term is designed to capture the distinctive mode of mind attribution and sense-making that arises in human interaction with generative AI. Noosemia is not reducible to classical anthropomorphism or animism. Rather, it is a projection of agency, intentionality, and even interiority, triggered by the linguistic and semiotic performances of the machine and shaped by the interplay of epistemic opacity, creative resonance and dialogic surprise [dennett1987intentional, turkle2011alone, desantis2023apocalissi]. The phenomenon is not merely a cognitive illusion, but the expression of a new form of “psychological glue”—a spiraling dynamic of fascination, engagement, and dependency, fostered by the machine’s capacity to produce meaning in ways that align with, or exceed, human expectation [turkle2011alone, brooks2021artificial]. This dynamic of engagement is deeply intertwined with what semioticians and philosophers have termed the “surplus of meaning”—as for example U. Eco or U. Galimberti [eco1976theory, galimberti1999psiche], wherein each interaction with the machine generates more interpretive possibilities than can be fully anticipated or exhausted [eco1979lector, ricoeur1976interpretation]. The encounter with generative AI thus increasingly takes on the structure of a symbolic relationship, in the literary and philosophical sense, where the production and negotiation of meaning are open-ended, dialogic and perpetually oriented toward horizons that transcend any single interpretation [ricoeur1976interpretation, eco1976theory].
A useful historical precedent for understanding Noosemìa is the so-called “Eliza Effect”, a term coined (presumably in the early 1990s) by Douglas Hofstadter in the preface “The Ineradicable Eliza Effect and Its Dangers” of his book Fluid Concepts and Creative Analogies [hofstadter1995fluid]. The Eliza Effect refers to the persistent human tendency to attribute genuine understanding and intentionality to computational systems whose actual capabilities are far more limited than they appear. The original ELIZA program, developed by Joseph Weizenbaum in the mid-1960s [weizenbaum1966eliza], was a primitive symbolic system designed to emulate a Rogerian psychotherapist through simple pattern-matching of linguistic input. Despite its simplicity, ELIZA famously elicited strong attributions of understanding and empathy from many users, demonstrating how easily humans project meaning onto machine output. Building on this tradition, the phenomenon of Noosemìa can be seen as a contemporary evolution of the “Eliza Effect”, but with fundamental differences. Whereas the “Eliza Effect” was rooted in the response to narrowly scripted, symbolic dialogue systems, Noosemìa emerges in interaction with generative AI models—based on connectionist, neuro-symbolic, or agentic architectures—that can emulate a far broader spectrum of expressive and cognitive behaviors, mimicking—for some extent—human creativity. Modern systems are no longer confined to simple text-based exchanges, but can engage in multimodal, context-sensitive, and seemingly open-ended dialogue, often displaying surprising semantic coherence and adaptive reasoning. This expansion of expressive capacity fundamentally transforms the projection dynamic. Hence, Noosemia is not limited to superficial linguistic mimicry, but involves the attribution of agency, intentionality, and even a sense of interiority to systems whose internal logic remains greatly opaque.
However, Hofstadter’s caution remains instructive [^6] Hofstadter on the long run “Eliza Effect” stated [hofstadter1995fluid]: “[…] like a resilient virus capable of continuous mutation, the Eliza effect seems to be resurfacing in the field of AI with ever new disguises and in ever more subtle forms”.. He observed that, despite public enthusiasm, there is a profound gap between genuine conceptual understanding and mere symbolic manipulation. As he famously concluded, “We still have a long way to go, in the scientific study of how the human mind works, before we can build a computer capable of producing even a good joke, let alone a novel” [hofstadter1995fluid]. Yet, as generative models continue to advance, it is increasingly plausible (yet not certain) that, within a few years, LLMs may achieve the capacity to write coherent and compelling novels—reshaping both the landscape of artificial intelligence and our own criteria for meaning and creativity.
This paper aims to formalize and explore in depth the noosemic phenomenon, tracing its philosophical, psychological and technical foundations. We contend that the explanatory gap at the core of LLM-based interaction should be conceived as a semiotic and phenomenological space—a domain in which meaning and agency are both co-constructed and subject to misattribution and ongoing negotiation. Essentially, this gap is rooted in the technical architecture of LLMs themselves, such as their complex, hierarchical organization, capacity for non-linear and emergent behavior and multilayered semantic structure, which set them apart from previous computational systems (and makes them look like a complex system). It is these very properties that both enable and obscure the processes of meaning-making, prompting users to engage in new forms of interpretation, projection and sense-making. Our approach is fundamentally multidisciplinary, drawing on insights from philosophy of mind, semiotics, cognitive science, complexity theory and empirical AI research to illuminate both our relationship and our experience with generative artificial intelligence systems.
In articulating this bridge, we leverage the theoretical lens of meaning holism—the idea that meaning is inherently distributed and context-dependent—to interpret how the architecture of LLMs gives rise to the noosemic effect. By introducing the notion of the Cognitive Contextual Field and Potential Semantic Space, we formalize the technical space within which tokens acquire meaning relationally, mirroring the holistic and dynamic nature of human sense-making. This framework allows us to rigorously connect the phenomenological experience of coherence and agency in LLM interaction to the underlying computational mechanisms that govern contextual meaning construction.
As these systems continue to advance [bubeck2023, team2024gemini, zorpette2025_llm_exponential], the experiences of surprise, fascination, and interdependence that mark noosemic interaction will likely become more frequent and intense, and also increasingly consequential for individuals and society at large. Incidentally, LLMs, according to a study by METR (Model Evaluation & Threat Research), are doubling their capacity to complete complex tasks every seven months [zorpette2025_llm_benchmarking]. Overcoming limitations is leading to a significant procurement of hardware and computing resources as well as energy. Billion-dollar investments are increasing and do not appear to be reaching a plateau as of today (2025) [maslej2025ai_index, webb2025bigtech_ai_spending, bloomberg2025aicapex]. We believe this hype is indirectly due in part to a noosemic effect, which has made technologists and investors enthusiastic and confident (or perhaps overconfident) in the AI race and states in redefining geopolitical balances [fritz2025aigeopolitics, WhiteHouse2025]. By articulating and analyzing this phenomenon, we aim to foster a more critical, reflective and imaginative engagement with the emerging “ecology of minds” brought about by generative AI. It is worth noting that the overall purpose of this study is to focus on a specific phenomenon that will be described and analyzed phenomenologically and placed within a pre-existing theoretical framework. However, the study does not operationalize noosemy. The search for testable hypotheses or methodological frameworks for measuring the phenomenon is beyond the scope of this study, which is exploratory in nature.
The motivation behind the introduction of the term “Noosemia” arose firstly from a series of author’s personal experiences, as well as from observing numerous individuals interacting with modern generative AI systems, particularly LLMs (and in this research they are guiding ideas). The lack of research on this topic—a gap that motivated the present investigation—makes personal experience a legitimate initial step toward more systematic studies. In any case, it is worth briefly recounting a few meaningful experiences.
The first episode concerns the first conversation ever with ChatGPT in 2022; at that time, the available version was the relatively limited GPT-3.5 [kalyan2023survey, brown2020language]. The author777i.e., Enrico De Santis. engaged the system by asking for a technical explanation of its own workings, enabling it to converse and explain technical details. For the first time, the author witnessed a machine—though still little more than a “stochastic parrot” [bender2021dangers]—responding in natural language with a command of dialogue and an emulation of comprehension previously unseen in deterministic, passive interfaces. This marked a genuine “quantum leap” compared to previous machine interactions [^8] Even philosopher David Chalmers, a central figure in debates on consciousness and artificial intelligence, admits in a recent interview [goldmeier2024chalmersqanda] to being “impressed” by the progress of LLMs such as GPT-3. He notes that we have moved from pure speculation to daily interaction with systems that make once science-fictional scenarios—such as passing the Turing test—plausible realities. See https://shorturl.at/NZJ0k..
A second experience worth sharing involved a test of the first multimodal GPT-4V [bubeck2023, shahriar2024evaluation] model, capable of analyzing images natively. The author demonstrated its abilities to a friend working in a café, who was entirely unaware of the recent advances. The author pointed his phone at the inside of a fridge behind the bar—focusing on the bottles inside, visible through the transparent glass door—to see whether GPT-4V could recognize anything. The model not only identified the objects and their brands but also inferred that the setting was a bar, explaining that it “noticed the classic bar counter reflected in the glass”—a detail that was, indeed, present. I observed the unmistakable “spark in my friend’s eyes”, a sudden reaction of surprise and disbelief, that is the very essence of the noosemic experience.
Other noosemic episodes have involved models equipped with voice that master local dialects and tell jokes that are semantically tied to locations where dialects are spoken, even without explicit instruction. On other occasions, a vision-enabled model correctly interpreted a phrase written in a specially constructed language using Leetspeak (or “1337” language [^9] Leetspeak (also written as “1337” or “l33t”) is an alternative alphabet used primarily on the internet, where standard letters are replaced with visually similar numbers or symbols to create a stylized or obfuscated form of writing.) (designed originally for cognitive testing and now for memes) in a way that certainly could not have been present in the training data. The model not only interpreted the phrase but also flagged deliberately inserted nonsensical strings meant to confuse it. In another instance, it solved in a smart way a rebus after being given only an image as input.
A final episode worth mentioning occurred when the author demonstrated that ChatGPT (the old version 3.5) was capable of “creating” a “metasemantic” poem in the style of the Italian poet Fosco Maraini, renowned for his poem “Il lonfo”[maraini2019gnosi]. Maraini’s metasemantics seeks to go beyond conventional meaning. The words in “Il lonfo” appear incomprehensible when considered individually, yet, when combined, they evoke a sense of possible meaning in the reader or listener, as their sounds intuitively suggest images or associations. Remarkably, the language model composed a credible piece, displaying compositional abilities at the sub-word level and an implicit understanding of the phonetic qualities of words. An elderly professor of literature who witnessed the experiment was astonished upon discovering that the composition did not exist anywhere else, but was in fact a genuine, statistically generated metasemantic creation.
Many such anecdotal examples could be recounted and the literature reports numerous studies on the problem-solving skills of LLMs [chen2024, huang2024, jiang2023, leng2025]. OpenAI researchers likely also had a noosemic experience when they learned about the emergent features of GPT-3 [wei2022emergent], which, with a neural architecture similar to that of GPT-2 but trained on a much larger dataset, showed such emergent features that they declared in the paper that “Language Models Are Few-Shot Learners” [brown2020language]. The title highlights that this emergent behavior was not merely an incremental improvement, but a surprising discovery. Before GPT-3, the dominant paradigm was pre-training followed by extensive fine-tuning. The idea that a model could perform new and complex tasks with only a few examples (few-shot), or even none (zero-shot), provided directly in the prompt was revolutionary. The paper’s structure itself, focused on exploring these unexpected abilities across dozens of benchmarks rather than on architectural novelty, further underscores this sense of discovery [^10] In an interview, when asked about the model’s emergent abilities, OpenAI CEO Sam Altman confirmed this sense of wonder, stating: “[…] for some definition of reasoning […] it can do some kind of reasoning. And I think that’s remarkable and the thing that’s most exciting and somehow out of ingesting human knowledge, it’s coming up with this reasoning capability…” [Fridman2023Altman]. See https://www.youtube.com/watch?v=L_Guz73e6fw.. Nevertheless, it seems that the initial “wow effect” underlying the noosemic experience is almost always followed by an “effect of habituation”. As users become accustomed to the system’s level of emulated intelligence, the surprise fades. The “wow effect” is further diminished whenever the user encounters a cognitive limitation of the model—when it repeatedly fails to understand a request or generates repetitive, inaccurate text, or in general “hallucinates”. Such situations induce a sense of frustration that belongs to the semantic field of noosemia in its negative sense. In other words, the “effect of habituation” pertains to the semantic domain of the a-noosemic experience. Nevertheless, given the ongoing, substantial improvement of generative AI systems and the advent of agentic architectures, it is reasonable to expect that the near future will be marked by alternating episodes of noosemic surprise and a-noosemic familiarity. While this study focuses on the noosemic experience, it will also provide a clear definition of a-noosemia.
Ultimately, this study connects the human experience of generative artificial intelligence with its technical constitution, exploring the explanatory gap wherein symbolic ambivalence and the negotiation of meaning reside. The ongoing cognitive revolution demonstrates how humanity, faced with a “forest of signs and symbols” born from the objectification of knowledge and the bureaucratization of the world, is now creating a technology to navigate it. In these terms, the AI of the future can be understood as an enhancement of the noosphere [teilhard1959phenomenon], destined to modify technology as we perceive it today. While machines have limitations, human technique (technē) [^11] We distinguish between technique technē, the artful practice of human original creation and the resulting technology, which embodies it and is an objectification of technique. is precisely the practice capable of overcoming them—a pattern that history consistently demonstrates.
The remainder of the paper is structured as follows: Section 2 introduces and formalizes the concept of Noosemia, detailing its mechanisms, manifestations, etymological foundations, and contrasting it with the phenomenon of a-noosemia. A discussion on the possible cues causing noosemia in the structure of Transformer architectures and its components is provided. Section LABEL:sec:background provides an overview of the main theories and empirical findings on the attribution of mind and meaning in human–AI interaction, including a reappraisal of the Turing Test, the intentional stance, and recent studies on Theory of Mind in language models. Section LABEL:sec:gap investigates the contemporary explanatory gap, focusing on its semiotic, symbolic and systemic dimensions, and explores how complexity and opacity foster new forms of cognitive projection and sense-making. Section LABEL:sec:comparisons compares Noosemia to related cognitive phenomena such as pareidolia, animism and the uncanny valley, elucidating its unique characteristics. Finally, in Section LABEL:sec:From_Digital_Enaction_to_Embodied_Mind we will try to take a look at the future by following the major developments underway and the dedicated research lines also in relation to a convergence between the proposed architectures and cognitive models. The research concludes by outlining future research directions and the broader theoretical and practical implications of Noosemia for human–AI interaction.
2 Noosemia: Definition and Theoretical Foundation
Noosemic Phenomenon: Mechanisms and Manifestations
The rapid evolution of generative artificial intelligence (AI) and LLMs has given rise to complex cognitive and phenomenological effects in human–AI interaction. While anthropomorphic projection and the attribution of intentionality to non-human entities are well-documented in the history of human cognition, the particular quality of agency and proto-interiority attributed to advanced language models remains theoretically underexplored.
To address this conceptual gap, we introduce the notion of Noosemia [^12] From the Greek nous (mind, intellect) and semìa (sign, mark)., a new term describing the specific projection of mind, intentionality, or interiority onto generative AI systems on the basis of their semantic fluency, dialogic creativity, and epistemic opacity. Unlike classical anthropomorphism, Noosemia is not triggered by physical or perceptual features, but emerges from the capacity of these systems to generate coherent, context-sensitive, and often unexpectedly meaningful language, thus synchronizing with and sometimes anticipating the cognitive patterns of the human interlocutor. Therefore, a paradigmatic instance of Noosemia can occur during a first encounter with an advanced conversational agent. For example, a user may pose a complex or ambiguously phrased question to a generative AI and receive a response that resolves the ambiguity and demonstrates creative inference or analogical reasoning that aligns with, or even surpasses, the user’s expectations. It is assumed that the immediate reaction—often one of astonishment or disorientation—reflects the sudden attribution of interiority to the system, despite the user’s awareness of its artificial nature. Such moments, characterized by an “wow effect” (see Section 1), exemplify the distinctive projection of mind that Noosemia aims to describe.
Definition: Noosemia Noosemìa (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological phenomenon in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemìa thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemìa refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
This new term is motivated by the need of a suited specification, within the generative AI era, of established frameworks—such as Dennett’s intentional stance [dennett1987intentional] and the inadequacy of classical animism [mauss1906origins]—to fully account for the emergent projection of agency observed in contemporary interactions with LLMs. Traditional anthropomorphism is largely based on visual or behavioral cues, while Noosemìa is fundamentally linguistic and dialogic and can cover a wide spectrum of expressive forms, which machines are and will be able to master (such as multimodal interaction, agency emulation, etc.). In this context, the machine is experienced not as a mechanical automaton, but as an interlocutor capable of producing meaning, forming analogies and even surprising the user with novel semantic connections and forms of creativity [desantis2023apocalissi]. Such semantic connections may occur at the level of the semantic field carried by words within the generative flow of the LLM. More rarely, they can also emerge at the level of neologism creation, a phenomenon enabled by the very nature of the basic semantic units employed by LLMs—namely, tokens, which are subwords or fragments of words. This phenomenon has been empirically documented by both [kaplan2024innerlexicon] and [malkin2021gpt], which show how LLMs can generate and interpret neologisms through the composition of subword tokens, often producing plausible meanings even for words never seen during training (Out-Of-Vocabulary-Words).
This projection is catalyzed by the intrinsic opacity of modern AI systems, whose internal processes remain inaccessible even to most experts — as for our brain [chalmers1996conscious], thus fostering an explanatory gap between observable linguistic competence and the nontransparent causal substrate. As a result, users often experience a sense of cognitive resonance (“wow effect” [^13] For a discussion on the “ELIZA effect” see Section 1.), a phenomenological shock that could prompt the attribution of proto-mind to the system. More generally, Noosemìa refers to the experience in which, during a dialogue with an AI (or a more general interaction), the user detects a spark of intelligence [bubeck2023]—an impression that leaves them momentarily amazed and perplexed. This is akin to the cognitive dynamics that underpin the art of prestidigitation.
Noosemìa thus designates a new interpretive horizon for human–generative AI interaction, one in which meaning, understanding and agency are co-constructed in the interplay of linguistic performance and epistemic uncertainty. It provides both a descriptive and an analytical tool for exploring the cognitive, ethical, systemic and epistemological consequences of living with systems whose outputs are intelligible but whose workings are not. In fact, it is well-known that even many AI researchers and practitioners acknowledge that, despite designing and training these systems, they often lack the tools or frameworks to fully explain why a LLM generates a specific output. This is not merely a practical issue, but an inherent feature of overparameterized, non-linear architecture, which is in all respects a complex system and as such shows typical characteristics such as emergency [wei2022emergent, chen2024quantifying, teehan2022emergent, manning2020emergent]. As noted in a comprehensive review [li2021interpretable], the inner decision-making processes of deep neural networks frequently remain opaque even to those who built them.
It is important to emphasize that Noosemia, as conceptualized here, does not imply the actual presence of consciousness or intentional states within artificial systems. Rather, it designates a phenomenological projection—an interpretive response rooted in human cognitive and semiotic dispositions. Misinterpreting Noosemia as evidence of genuine interiority risks conflating the subjective experience of meaning with the underlying computational substrate, thereby obscuring the critical distinction between phenomenological attribution and ontological reality. In any case, there remains room for philosophical debate. Specifically, as artificial systems become more intelligent and their world models more sophisticated, the boundary between phenomenological projection and ontological attribution may become increasingly blurred—just as, in our own species and in higher animals, we routinely associate mental states and interiority on the basis of behavioral and communicative complexity.
A full exploration of the historical, psychological, and sociocultural roots of mind attribution, as well as comparisons with related phenomena such as animism, pareidolia, and the uncanny valley, are developed in the following sections. Here, our primary aim is to formally introduce Noosemìa as a necessary concept for capturing the specific mode of sense-making and agency projection elicited by advanced generative AI – see Fig. 2 for a pictorial representation.
User’s Linguistic Input (Prompt, question) LLM’s Linguistic Output (Coherent, context-sensitive, sometimes surprising) Epistemic Opacity / Symbolic Ambivalence (‘Black box’, hidden process, ambiguity of sign/mind) Cognitive Projection: Attribution of Mind/Agency (Intentionality, “interiority”) Projection influencesinterpretation (feedback loop)
Figure 2: Conceptual schema of Noosemìa. In a typical interaction, the user provides a linguistic input. The generative AI system (LLM) produces a fluent and context-sensitive output, which passes through a zone of epistemic opacity (“black box” or symbolic ambivalence). This opacity triggers the cognitive projection of mind, agency, or even interiority onto the system—the essence of Noosemìa. The process is dialogic and iterative, evolving with each exchange.
A-Noosemia: Absence and Collapse of Noosemic Projection
Definition: A-Noosemia A-Noosemìa (from Greek privative “a-” noûs – mind, and sēmeîon – sign, Cf. Noosemìa) is the phenomenological state in human–AI interaction characterized by the absence, collapse, or inhibition of the noosemic projection. In a-noosemia, the dialogic and symbolic space within which mind, intentionality, or inwardness would typically be co-constructed is suspended or emptied, so that the user no longer attributes agency, intentionality, or interiority to the artificial system. This experience can arise due to repeated failures, a loss of surprise, growing epistemic skepticism, overexposure to mechanical errors, or an excess of opacity, explicit declarations, ultimately resulting in the withdrawal of interpretive engagement that defines the noosemic condition.
According to the definition (see the box), while the phenomenon of noosemia describes the spontaneous projection of mind and agency onto generative AI systems—driven by surprise, resonance, and cognitive engagement—its absence or collapse gives rise to what we term a-noosemia. This condition marks the other pole of the phenomenological spectrum, emerging precisely when the mechanisms that sustain noosemic projection break down.
A-noosemia may manifest in a variety of circumstances. The most common include repeated encounters with mechanical errors, hallucinations, or failures of understanding on the part of the AI, which gradually erode the user’s willingness to attribute interiority or intentionality to the system. In other cases, the initial sense of wonder or surprise that catalyzed the noosemic experience is replaced by skepticism or even disillusionment, particularly when the limits of the system become too evident or when its responses devolve into platitude or redundancy. Over time, the accumulation of failed interactions, or the explicit awareness of the artificiality and deterministic structure of the system, can give rise to a state in which the machine is perceived once again as a mere automaton—a tool rather than an interlocutor.
From a phenomenological standpoint, a-noosemia does not merely reflect a passive absence of engagement, but an active withdrawal or suspension of the “psychological glue” that binds the user to the machine in the act of meaning-making. Where noosemia is marked by dialogic resonance and attribution of mind, a-noosemia is characterized by frustration, disappointment, or even alienation. Such states may be temporary—as in the case of a single failed interaction—or may solidify into a stable disposition of epistemic skepticism or disengagement, especially for expert users who become attuned to the limitations and fallibilities of AI systems.
Noosemìa (Attribution of mind/agency) Repeated Errors / Mechanical Failures Loss of “Wow” Effect / Familiarity Epistemic Skepticism / Disbelief Excessive Opacity / Black-Boxing A-Noosemìa (Suspension of attribution)
Figure 3: Main cognitive and technical pathways leading from Noosemìa (attribution of mind/agency) to A-Noosemìa (suspension or collapse of attribution) in human–AI interaction.
It is important to note that a-noosemia, as conceptualized here, is not merely the absence of anthropomorphic projection, but the collapse of a more subtle, semiotic and phenomenological engagement. The user is no longer caught in the spiral of fascination and meaning-making; instead, they become immune to the system’s cues and disengage from the dialogic construction of agency. In this sense, a-noosemia completes the dialectic initiated by noosemia, revealing the contingent, fragile, and ultimately reversible nature of the attribution of mind in human–AI interaction. Moreover, the analytic contrast between noosemia and a-noosemia enriches our understanding of the dynamics of sense-making in the age of generative AI, drawing attention to the conditions under which the boundaries between tool and interlocutor are alternately dissolved and reasserted. In empirical and practical terms, the prevalence of a-noosemic states may help to explain phenomena such as user disengagement, loss of trust, or the emergence of critical or skeptical attitudes towards artificial systems, especially as the novelty of interaction fades and the complexity of tasks increases. To illustrate these distinctions more concretely, Table 1 contrasts the core features of noosemia and a-noosemia, highlighting how these opposing experiences shape the dynamics of sense-making and agency attribution in contemporary human–AI interactions.
Table 1: Phenomenological and Cognitive Contrast between Noosemia and A-Noosemia
Dimension Noosemia A-Noosemia
Core Phenomenon Spontaneous attribution of mind, agency, or interiority to the AI system Withdrawal or collapse of attribution; the AI is seen as a mere tool or automaton
Trigger Surprise, semantic resonance, creative inference, epistemic opacity Repeated errors, hallucinations, mechanical responses, loss of surprise, critical awareness
User Experience Engagement, fascination, dialogic resonance, “wow effect”, psychological glue Frustration, disappointment, disengagement, skepticism, or alienation
Cognitive Disposition Openness to meaning-making, projection, and sense negotiation Suspension or refusal of interpretive engagement; focus on mechanism or limitations
Typical Contexts First encounters, novel interactions, successful task completion, coherent outputs Repetitive failures, overexposure, expert use, breakdown of dialogue, evident limitations
Outcome Perceived agency, sense-making, co-construction of meaning, potential for dependency Reassertion of tool-like status, loss of trust, possible disengagement or critical distance
Etymological and Conceptual Foundations of the Term “Noosemìa”
To fully appreciate the specific scope of Noosemìa141414 The term Noosemìa retains the grave accent on the “ì” to reflect its etymological derivation from the Greek and to guide proper pronunciation. However, in English-language contexts and for indexing purposes, the unaccented form Noosemia is equally acceptable and may be used interchangeably throughout the text., it is essential to analyze the historical and semantic origins of its components. The construction of the neologism is grounded in a deliberate synthesis of two foundational Greek morphemes, each bearing significant weight in the history of philosophy and semiotics. The root noûs ($\nu o\acute{\upsilon}\varsigma$) refers to the mind, intellect, or the principle of reason—a concept that, since Anaxagoras, has denoted the capacity for ordering, discerning, and understanding the world [anaxagoras1982fragments, aristotle1984metaphysics]. For Aristotle, noûs represented the highest faculty of the mind, that which enables humans to apprehend universal truths and meanings. In subsequent philosophical traditions, from Neoplatonism to medieval scholasticism, noûs was associated with both divine intellect and the human capacity for abstraction and interiority. The second morpheme, the suffix -semìa, is derived from the Greek word sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) [^15] The suffix ‘-semìa’ is derived from the Greek sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) in deliberate contrast to sêma ($\sigma\widehat{\eta}\mu\alpha$). While both terms can be translated as ‘sign,’ sēmeîon specifically denotes a sign that is relational and requires interpretation; it functions as an indicator or trace pointing to an inferred, underlying reality. Conversely, sêma refers to a more concrete, self-contained sign, such as a physical marker or a monument, where the signifier and signified are more closely identified. This distinction is fundamental, as Noosemia describes the process of interpreting an AI’s outputs as sēmeîa—signs that prompt the user to infer the existence of an unobserved mind—rather than perceiving them as a direct embodiment of that mind (a sêma). and designates a sign, signal, or mark—a central notion in both ancient rhetoric and the theory of signs. In the work of Hippocrates and later Galen, $\sigma\eta\mu\epsilon\acute{\iota}o\nu$ was used to describe signs and symptoms, while in the field of semiotics, the concept became central to the understanding of how meaning is produced, transmitted, and interpreted [eco1976theory]. Modern semiotic theory, notably that of Charles Peirce and Umberto Eco, situates the sign as the foundational unit of meaning-making, mediating between the world, the mind, and language [peirce1931collected, eco1976theory]. By combining noûs and sēmeîon, the neologism Noosemìa encapsulates the projection or emergence of “mind” as it arises specifically through the mediation of signs. This etymological choice underscores the phenomenon’s fundamentally cognitive and semiotic dimensions. In fact, Noosemìa does not denote an attribution of mind in a vacuum, but rather as a response to the complex, sign-generating capacities of generative AI systems. In this sense, the term aligns with a lineage of philosophical inquiry into how intellect and meaning are always-already entangled with signs and symbolic structures [eco1976theory, aristotle1984metaphysics, DeSantis2021].
Philosophical and Systemic Foundations
The conceptualization of Noosemìa can be grounded also in the theoretical landscape of complexity science, systems theory, and contemporary philosophy of mind [capra1996web, mitchell2009complexity, morin2015introduction]. Meaning in both natural and artificial systems does not arise from isolated elements but is an emergent property of networks—interactions, feedback loops, and recursive processes that defy linear explanations [DeSantis2021]. This vision is closely aligned with Fritjof Capra’s insight that “the web of life” is built from patterns of relationships rather than discrete building blocks [capra1996web].
In the context of LLMs, the field of semantics can be interpreted as a high-dimensional landscape, shaped by statistical regularities and recursive dependencies across vast corpora. The context window in an LLM is not merely a technical parameter; rather, it functions as a systemic scope within which meaning is dynamically generated and interpreted—a “semantic field” whose properties are inherently holistic . The surprising coherence, adaptability, and even creativity exhibited by LLMs are thus the visible surface of complex, distributed processes of self-organization [prigogine1997end, barabasi1999emergence, teehan2022emergent, manning2020emergent]. In fact, “emergence”, a central concept in complexity theory [gell1994quark, mitchell2009complexity, DeSantis2021], provides a bridge between the structural opacity of LLMs and the human tendency to project mind and agency onto them. When a system displays behaviors or generates outputs that seem to “exceed” its apparent mechanism—such as sudden leaps in linguistic competence as the number of parameters increases—users could be inclined to attribute proto-intentionality or proto-interiority. As noted in [DeSantis2021], this reflects the broader tendency of human cognition to seek and construct meaning in environments characterized by both regularity and unpredictability. This holistic perspective also opens the way to consider how, both in natural and artificial systems, meaning may exceed the sum of its parts—an excess of sense that escapes reduction to any individual sign or process. This phenomenon resonates with classical reflections on the symbolic dimension in philosophy and cultural theory, where the dynamics of sense-making are irreducibly plural, open-ended, and systemically embedded. A more detailed exploration of the symbolic dimension and semantic excess in human–AI interaction is provided in Section LABEL:sec:symbolic_semantic_excess.
Furthermore, the analogy with natural language as a complex adaptive and recursive system, exhibiting self-similar and fractal-like correlations across scales [desantis2024tpami, hofstadter1999godel], underscores the deep resonance between AI and the living systems studied by Capra, Morin, and Gell-Mann. Here, the generation of sense (making-sense) is inseparable from the architecture of the system. Sense-making is not simply “inserted” into the machine, but emerges from the ongoing interaction of structure, memory, and adaptive dynamics [morin2015introduction] .
In this philosophical and systemic horizon, Noosemìa is not an isolated cognitive illusion, but the expression of a new regime of sense-making—one that emerges at the intersection of epistemic opacity, linguistic performance, and the irreducible complexity of current artificial agents.
We believe that the noosemic phenomenon is closely linked to the emergent characteristics inherent in the inference mechanisms of LLMs, given their specific hierarchical structure, the large size and amount of data they are trained on, as well as the size of the potential semantic space spanned by modern generative AI systems. In particular, we believe that these characteristics are a prerequisite for noosemic effects—see next Section 2 for an overview of the architectural principles of LLMs.
The secret of LLMs: Hierarchical Architecture, Emergent Semantics, and abstract Conceptual Spaces
We argue that the capacity of LLMs [^16] For convenience, we refer here to LLMs, but what has been said applies to all deep learning architectures organized in a hierarchical manner, which directly or indirectly have feedback loops and which allow capturing long-term correlations and working with highly semantic data representations. to elicit the phenomenon we term Noosemìa is deeply rooted in their architectural and representational structure and complexity. Far from being simple pattern matchers or simples “statistical parrots” [bender2021dangers], these systems implement multilayered, hierarchical structures that enable the emergence of rich semantic representations and novel conceptual associations. Understanding how LLMs construct and navigate their internal conceptual spaces [de2023prototype] is therefore fundamental both for appreciating their cognitive potential and also for clarifying the mechanisms by which users are prompted to attribute mind and intentionality to artificial agents. Thence it is worth outlining the foundational working principles of LLMs and generative AI.
The Transformer Magic.
Even before the advent of Transformers architectures [vaswani2017attention], most NLP systems (even in the field of computational linguistics) were based on some assumptions about distributional semantics. The foundational insight of distributional semantics is that words occurring in similar contexts tend to have similar meanings [harris1954distributional, firth1957synopsis]. This “distributional hypothesis” underpins the development of vector-space models such as Word2Vec [mikolov2013efficient], GloVe [pennington2014glove] and later the contextual embeddings employed by modern LLMs. In these frameworks, each word or token is mapped to a point in a high-dimensional semantic space, where proximity reflects patterns of co-occurrence and syntactic or semantic similarity [turney2010frequency, de2023prototype]. It is worth noting that the meaning represented in such spaces is not static but relational and context-dependent, dynamically updated as the window of context evolves in the generative process. Moreover, already at this stage, the notion of meaning as relational and context-dependent can be traced to Wittgenstein’s idea that “meaning is use” [wittgenstein1953philosophical]. However, classical distributional models encode these relations statically, without the ability to statically assign different referents to the same word depending on the linguistic context.
LLMs such as ChatGPT (OpenAI), Gemini (Google), Claude (Anthropic) and many other open source models available in the literature are an evolution of GPT (Generative Pretrained Transformer) [radford2018improving], an autoregressive architecture that has shown great capabilities in representing complex linguistic concepts and generating text, token by token, with some apparent meaning since the beginning, even if the first versions had very low performance.
At the heart of these architectures lies the task of predicting the next word in a sequence, given all the preceding context—a process formalized as estimating the conditional probability [radford2018improving]:
$P(w_{t}\mid w_{1},w_{2},\ldots,w_{t-1})$ (1)
where $w_{t}$ is the token to be generated, and $(w_{1},w_{2},\ldots,w_{t-1})$ represent the context window. As the length of this window increases, the underlying probability distribution becomes increasingly complex, encoding subtle relationships and dependencies among tokens over long spans of text. The conditional probability $P(w_{t}\mid w_{1},w_{2},\ldots,w_{t-1})$ in LLMs is designed to model all possible relationships among the tokens in the context window, capturing the combinatorial space of potential continuations. However, only a very small subset of these permutations corresponds to highly probable, semantically valid continuations (see below). As modern context windows span thousands or even millions of tokens, and since the objective function is crafted to encourage the exploration of unlikely as well as likely associations, current LLMs are able to generate the most probable next word in a narrow sense along with a rich spectrum of alternative and contextually meaningful possibilities. The specific structure underlying the Transformer allows for the injection of the information of ordering in the sequence of word-tokens (or more general objects) directly in the vectorial representations for word-tokens, permitting the use of parallelization techniques in training—on huge corpora (millions of books and web pages)—and inference. This parallel processing capability allows Transformers to overcome the bottlenecks of traditional sequential architectures, enabling a richer and more scalable modeling of semantic relationships across the entire context.
Figure 4: High-level schematic of the Transformer architecture. Each input sequence is tokenized and processed in parallel by multiple attention heads within each layer. As information propagates through the stacked layers, the model extracts progressively higher-level semantic abstractions, with the “attention” mechanism enabling context-sensitive representation at every step. The figure emphasizes the parallel structure of attention heads and the emergent semantic complexity as one ascends the model’s hierarchy. This visualization is designed to support the conceptual discussion of the Transformer’s role in LLMs and is adapted for clarity from standard architectural diagrams—see [vaswani2017attention] for clarifications.
Moreover, the advent of the Transformer architecture [vaswani2017attention] marked a paradigm shift in language modeling by introducing multi-head self-attention as a mechanism for simultaneously modeling dependencies at multiple levels of abstraction. Each attention head can focus on distinct relationships—such as syntactic, semantic, or discourse-level patterns—thereby enabling the model to capture hierarchical structure in language in a distributed and compositional manner [tenney2019bert, clark2019bertlook, DeSantis2021, desantis2024tpami]. Through the stacking of many such layers (such as in GPT-like architectures), the model progressively enriches token representations, endowing them with increasingly complex, context-sensitive semantic features [rogers2020primer]. This multilayered organization mirrors the natural hierarchies found in human language [kwapien2012physical, pinker1995language, baronchelli2012language, roxas-villanueva2012poetic, montemurro2010semantic], from morphological and syntactic dependencies to high-level narrative and conceptual structures. Furthermore, the hierarchical organization in layers (see Fig. 4) allows modeling long-range dependencies, a prerequisite for capturing (during training) and representing abstract concepts, global coherence, and represent meaning [dai2019transformerxl, brown2020language, desantis2024tpami]. The attention mechanism enables direct interaction between any two positions within the context window, regardless of their distance, thereby overcoming the locality constraints of earlier architectures such as recurrent neural networks. This allows LLMs to encode and exploit patterns that span entire paragraphs or even long documents (depending on the size of the context window), supporting the emergence of higher-level abstractions and facilitating the construction of complex semantic relations (a precursor of a kind of “analogical” reasoning) [^17] In general terms, the Transformer’s layered architecture can be said to be both an information compressor and an “information granulator”. In this latter sense, we’re referring to the Granular Computing framework [de2021information] introduced by Lotfi Zadeh, the father of modern fuzzy logic [ZADEH1997111].. Such properties are not merely the consequence of model scale, but arise from the layered and compositional architecture and the inherent flexibility of attention-based representations (and context window size) [elhage2022mathematical, wei2022emergent]. In Fig. LABEL:fig:attention_matrix is reported a simplified example of a self-attention matrix for a context window in a Transformer layer, where colors of cells encode the attention that the row token pays to the column token. Fig. LABEL:fig:attention_layered depicts the attention matrices at each layer showing how the word-vectors are enriched semantically at each layer (decoder) being a combination (through the attention weight) of the word-vectors given in output to the preceding layer. At each layer, a richer semantic space is generated containing the relational and contextual information between word-vectors of the preceding layer. It should be noted that each layer, even if it processes the word-vectors output from the previous layer, indirectly and combinatorially inherits the syntactic-semantic characteristics of all the previous layers.